# Cross-positive linear maps, positive polynomials and sums of squares

### v1

@@ 4. Blekhermann type volume estimates
$\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ ::=:= $\left\{p\in\mathcal{Q}\colon p(x,y)\geq 0\quad\text{for all }(x,y)\in V(I)\right\},$
$\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ ::=:= $\left\{p\in\mathcal{Q}\colon p-\sum_{i=1}^{k}p_{i}^{2}\in I\quad\text{for some }k\in\mathbb{N}\text{ and }p_{i}\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}\right\},$
@@
We will estimate the gap between the cones $\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ and $\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ by comparing the volumes of suitably chosen compact sections of these cones. First we have to carefully introduce an appropriate measure on the set $(\mathbb{S}^{n-1}\times\mathbb{S}^{n-1})\cap V(I)$ w.r.t. which we will integrate elements from $\mathcal{Q}.$ This is the content of the next subsection.
@@ 4.1. Definition of integration
and equip it with the subspace topology. Let $C(T)$ denote the vector space of all continuous functions on $T$. The special orthogonal group $\operatorname{SO}(n)$ acts on the vector space $C(T)$ by rotating the coordinates, i.e., for $g\in\operatorname{SO}(n)$ and $f\in C(T)$ we have
@@
is homeomorphic to $\operatorname{SO}(n-2).$ We denote by $\hat{\phi}_{w}:\operatorname{SO}(n)/\ker(\phi_{w})\to T$ the induced map of $\phi_{w}$ on $\operatorname{SO}(n)/\ker(\phi_{w})$, which is the set of left cosets of $\ker(\phi_{w})$ in $\operatorname{SO}(n)$. Let $\widehat{\sigma}$ be the Haar measure on $\operatorname{SO}(n)$. We equip the quotient space $\operatorname{SO}(n)/\ker(\phi_{w})$ with the positive normalized $\operatorname{SO}(n)$-invariant measure $\sigma$ induced by $\widehat{\sigma}$ which exists and is unique. (See [Nac65, Theorem 1 on p. 138] and use the fact that compact groups are unimodular for uniqueness.)
@@
$(\hat{\phi}_{w})^{-1}(g\Delta)$ =$=\{g^{\prime}\in\operatorname{SO}(n)/\ker(\phi_{w})\colon g^{\prime}w\in g\Delta\}=\{g^{\prime}\in\operatorname{SO}(n)/\ker(\phi_{w})\colon g^{-1}g^{\prime}w\in\Delta\}$
@@
We already established the existence of a measure in Proposition 4.1. It remains to prove the uniqueness. Let us assume on the contrary that $\mu_{1}$ and $\mu_{2}$ are two different normalized $\operatorname{SO}(n)$-invariant measures on $T$. Then $(\phi^{-1}_{w})_{\ast}(\mu_{1})$ and $(\phi^{-1}_{w})_{\ast}(\mu_{2})$ are two different normalized $\operatorname{SO}(n)$-invariant measures on $\operatorname{SO}(n)/\ker(\phi_{w})$. But this contradicts the uniqueness of $\sigma$. ∎
@@
In literature [Mui82, Chi03] the set $T$ is known as the Stiefel manifold $V_{2,n}(\mathbb{R})$ of all $2$–frames in $\mathbb{R}^{n}$, i.e., the sets of all pairs of orthonormal vectors in $\mathbb{R}^{n}$. Equivalently, $V_{2,n}(\mathbb{R})$ is the set of all real $n\times 2$ matrices $X$ such that $X^{T}X$ is the $2\times 2$ identity matrix. Regarding $T$ as a manifold, it can also be equipped with the uniform normalized measure w.r.t. the action of the orthogonal group $\operatorname{SO}(n)$ [Chi03, §1.4.3]. By the discussion above, this measure coincides with the measure $\sigma$ introduced above.
@@ 4.2. Estimates
Now that we defined the measure on $T$ we can construct the appropriate sections of $\operatorname{Pos}_{\mathcal{Q}}$ and $\operatorname{Sos}_{\mathcal{Q}}$ and present the volume estimates for these sections (see Theorems 4.5, 4.6 below).
@@
$\begin{split}\widetilde{\operatorname{Pos}}^{(n)}_{\mathcal{Q}}&=\left\{f\in\mathcal{Q}\colon f+(\sum_{i=1}^{n}x_{i}^{2})(\sum_{i=1}^{n}y_{i}^{2})\in\left(\operatorname{Pos}^{(n)}_{\mathcal{Q}}\right)^{\prime}\right\},\\ \widetilde{\operatorname{Sos}}^{(n)}_{\mathcal{Q}}&=\left\{f\in\mathcal{Q}\colon f+(\sum_{i=1}^{n}x_{i}^{2})(\sum_{i=1}^{n}y_{i}^{2})\in\left(\operatorname{Sos}^{(n)}_{\mathcal{Q}}\right)^{\prime}\right\}.\end{split}$
@@
(4.2) =$\mathcal{M}=\left\{f\in\mathcal{Q}\colon\int_{T}f\;{\rm d}\sigma=0\right\}.$
Since $\sigma$ is normalized, we have that
@@
The proof of Lemma 4.4 is the same as the proof of [KMŠZ19, Lemma 1.4].
We are now ready to compare the volumes of the sections defined above. The proofs, similar to those in [KMŠZ19], are given in Appendices A and B. The lower bound for the volume of the section of nonnegative biforms from $\mathcal{Q}$ is as follows:
@@
Proof.
See Appendix B. ∎
Remark 4.7.
If we want to compare the sizes of two cones $K\subseteq L\subseteq\mathbb{R}^{n}$ in a fixed metric, the most unbiased choice of a compact set $C$ to compare the sizes of $K\cap C$ and $L\cap C$, is the unit ball $B$ of this metric. In our case, the metric is the $L^{2}$ norm, coming from the inner product (4.3). In this norm, the condition $f\in B$ is given by a quadratic inequality in the coefficients of $f$ and therefore sharp lower and upper bounds on $K\cap B$ (resp. $L\cap B$) following the same asymptotics are difficult to establish. The idea of replacing the unit ball $B$ with a hyperplane whose normal is some vector from the unit sphere leads to more manageable conditions. The choice of the hyperplane is not arbitrary, since its position can have a large effect on the size difference of the intersections, e.g., if the normal is almost perpendicular to some ray on the boundary of the largest cone $L$, the difference in size can be very large, even if the smaller cone is not significantly smaller. However, if the cones have a common line of symmetry, then the hyperplane, whose normal is this line of symmetry, fairly captures size difference between the cones. In our case, the polynomial $(\sum_{i=1}^{n}x_{i}^{2})(\sum_{i=1}^{n}y_{i}^{2})$ is a fixed point for the action of the orthogonal group $O(n)$ on $\mathbb{R}[{\tt x},{\tt y}]_{2,2}$, defined by $O\cdot p({\tt x},{\tt y})=p(O^{-1}{\tt x},O^{-1}{\tt y})$. Note that the ideal $I$ is invariant under this action and therefore the action extends naturally to the action on $\mathcal{Q}$. It is clear that the sets $\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ and $\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ are invariant for this action and therefore comparing their sizes by intersecting them with $\mathcal{H}_{\mathcal{Q}}^{(n)}$ is an appropriate choice.
Blekherman [Ble06, Theorem 6.1] established volume bounds for sum of squares forms. Our proofs of Theorems 4.5 and 4.6 freely borrows from his ideas. An important ingredient in the proof of Theorem 4.6 is also a new version of the Reverse Hölder inequality, which we prove in Subsection 4.3 below.
@@
Remark 4.9.
In [Ble06] Blekherman proved that for a fixed degree bigger than 2 the ratio between the volume radii of compact sections of the cones of sum of squares forms and nonnegative forms goes to 0, as the number of variables goes to infinity. Corollary 4.8 is an analog of his result for sum of squares biquadratic forms and nonnegative biquadratic forms on Stiefel manifolds $V_{2,n}(\mathbb{R}).$ (See Remark 4.3.)
@@
For every $n\in\mathbb{N}$ the probability $p_{n}$ that a cross-positive map $\Phi:M_{n}(\mathbb{R})_{\mathrm{sym}}\to M_{n}(\mathbb{R})_{\mathrm{sym}}$ is completely cross-positive, is bounded above by
@@ 4.3. Reverse Hölder inequality
We write $I_{1,1}=I\cap\mathbb{R}[{\tt x},{\tt y}]_{1,1}$. A bilinear form $g\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}/I_{1,1}$ is called symmetric (resp. skew-symmetric) if it of the form $g({\tt x},{\tt y})={\tt x}^{T}A{\tt y}+I_{1,1}$ for some symmetric (resp. skew-symmetric) matrix $A\in\mathbb{R}^{n\times n}$.
@@
For a bilinear biform $g\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}/I_{1,1}$ we have
@@
If $g$ is symmetric, then we can take $\sqrt{3}$ instead of $\sqrt{6}$ in (4.4) above, while if $g$ is skew-symmetric, $\sqrt{6}$ can be replaced by $4]{6}$. Moreover, the constants $\sqrt{3}$ (resp. $4]{6}$) are asimptotically, as $n\to\infty$, sharp.
@@
The proof of Proposition 4.11 will be done separately for the symmetric (Subsection 4.3.2) and skew-symmetric case (Subsection 4.3.3), while the general case (Subsection 4.3.4) follows by the fact that every bilinear form $g$ can be written as a sum of a symmetric form $g_{s}$ and a skew-symmetric form $g_{a}$, together with the observation that $g_{s}$ is perpendicular to $g_{a}$ in the $L^{2}$ inner product. For the proof we first need to compute the values of the integrals of monomials of bidegree $(2,2)$ and some monomials of bidegree $(4,4)$ with respect to $\sigma$, which is the content of Subsection 4.3.1. Using these computations, (4.4) becomes an inequality in the coefficients of $g_{a}$ (resp. $g_{s}$), which holds.
@@
In [Duo87] a version of the Reverse Hölder inequality with respect to the Lebesgue measure on the unit sphere and polynomials of any degree is established. [KMŠZ19, Lemma 2.9] extends this result to the product measure of two Lebesgue measures on unit spheres. However, in the proof of Theorem 4.6 we cannot use this extension, because the measure $\sigma$ is not a product measure. Therefore we have to establish the Reverse Hölder inequality we need in our setting.
@@
$I_{6}$ $=\int_{T}{\tt z}_{ij}{\tt z}_{kl}\;{\rm d}\sigma=0\quad\text{if at least one of }i,j,k,l\text{ occurs an odd number of times},$
$I_{7}$ $=\int_{T}{\tt v}_{ij}{\tt v}_{kl}\;{\rm d}\sigma=0\quad\text{if at least one of }i,j,k,l\text{ occurs an odd number of times,}$,
$I_{8}$ $=\int_{T}{\tt w}_{ij}{\tt w}_{kl}\;{\rm d}\sigma=0\quad\text{if at least one of }i,j,k,l\text{ occurs an odd number of times,}$,
@@
for $\text{for }i,j,k,l=1,\ldots,n,\;i,j,k\text{ pairwise different},$
@@
for $\text{for }n\geq 4\text{ and }i,j,k=1,\ldots,n,\;i,j,k,l\text{ pairwise different},$
@@
for $\text{for }n\geq 4\text{ and }i,j,k,l\text{ pairwise different},$
@@
if at least one of $\text{if at least one of }i,j,k,l,o,p,r,s\text{ occurs an odd number of times,}$,
@@
if at least one of $\text{if at least one of }i,j,k,l,o,p,r,s\text{ occurs an odd number of times.}$.
@@
Then $H_{n}^{1}(\underline{\phi})$ is (e.g., the formula for $L_{1}(\theta)$ in [Tum65, p. 3–4] or use induction on $n$)
@@
The set $T=V_{2,n}(\mathbb{R})$ (see Remark 4.3) can be parametrized by (see, e.g., [Chi03, p. 48–49] or [Tum65, §2])
@@
where by [Chi90, Theorem 2.1] (taking $V=x(\underline{\phi})$, $G(V)=H_{n}^{1}(\underline{\phi})\text{ without the first column}$ and $Z=\text{the first column of }H_{n-1}^{1}(\underline{\psi})$) and [Blu60],
@@
By the invariance of the integral with respect to the change of indices we can assume without loss of generality that $i,j,k,l\in\{1,2,3,4\}$ in all equalities of Lemma 4.13. Due to the difference in parametrizations of some of the coordinates $x_{i}(\underline{\phi})$, $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3,4$, we separate cases $n\geq 6$, $n=5$, $n=4$ and $n=3$ in the rest of the proof.
Case 1: $n\geq 6$. The coordinates $x_{i}(\underline{\phi})$, $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3,4$, are equal to:
@@
The fact $I_{3}=\frac{n+1}{n-1}I_{1}$ follows from the following computation:
11 11 $=\int_{T}\Big{(}\sum_{i=1}^{n}{\tt x}_{i}^{2}\Big{)}\Big{(}\sum_{i=1}^{n}{\tt y}_{i}^{2}\Big{)}\;{\rm d}\sigma=\int_{T}\Big{(}\sum_{i=1}^{n}{\tt z}_{i}^{2}\Big{)}\;{\rm d}\sigma+\int_{T}\Big{(}\sum_{i\neq j}{\tt z}_{ij}^{2}\Big{)}\;{\rm d}\sigma=nI_{1}+n(n-1)I_{3}.$
@@
Next we prove that $\int_{T}{\tt z}_{ij}{\tt z}_{kl}\;{\rm d}\sigma=0$ if at least one of $i,j,k,l$ occurs an odd number of times. Write $g({\tt x},{\tt y}):={\tt x}_{i}{\tt y}_{j}{\tt x}_{k}{\tt y}_{l}$ and let $i_{1}$ be the index among $i,j,k,l$, which occurs an odd number of times. Since
@@
Case 2: $n=5$. Note that the parametrizations of $x_{i}(\underline{\phi})$, $i=1,2,3,4$, and $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3$, are the same as in the case $n\geq 6$, while
@@
00 0 $=\int_{T}({\tt z}_{1}+{\tt z}_{2}+{\tt z}_{3}+{\tt z}_{4}+{\tt z}_{5})^{4}\;{\rm d}\sigma=\sum_{i=1}^{5}\int_{T}{\tt z}_{i}^{4}\;{\rm d}\sigma+4\sum_{i\neq j}\int_{T}{\tt z}_{i}^{3}{\tt z}_{j}\;{\rm d}\sigma+6\sum_{i<j}\int_{T}{\tt z}_{i}^{2}{\tt z}_{j}^{2}\;{\rm d}\sigma$

### v2

@@ 4. Blekherman type volume estimates
$\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ $:=$ $\left\{p\in\mathcal{Q}\colon p(x,y)\geq 0\quad\text{for all }(x,y)\in V(I)\right\},$
$\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ $:=$ $\left\{p\in\mathcal{Q}\colon p-\sum_{i=1}^{k}p_{i}^{2}\in I\quad\text{for some }k\in\mathbb{N}\text{ and }p_{i}\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}\right\},$
@@
We will estimate the gap between the cones $\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ and $\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ by comparing the volumes of suitably chosen compact sections of these cones. First we have to carefully introduce an appropriate measure on the set $(\mathbb{S}^{n-1}\times\mathbb{S}^{n-1})\cap V(I)$ with respect to which we will integrate elements from $\mathcal{Q}.$ This is the content of the next subsection.
@@ 4.1. Definition of integration
and equip it with the subspace topology. Let $C(T)$ denote the vector space of all continuous functions on $T$. The special orthogonal group $\operatorname{SO}(n)$ acts on the vector space $C(T)$ by rotating the coordinates, i.e., for $g\in\operatorname{SO}(n)$ and $f\in C(T)$,
@@
is homeomorphic to $\operatorname{SO}(n-2).$ We denote by $\hat{\phi}_{w}:\operatorname{SO}(n)/\ker(\phi_{w})\to T$ the induced map of $\phi_{w}$ on $\operatorname{SO}(n)/\ker(\phi_{w})$, which is the set of left cosets of $\ker(\phi_{w})$ in $\operatorname{SO}(n)$. Let $\widehat{\sigma}$ be the Haar measure on $\operatorname{SO}(n)$. We equip the quotient space $\operatorname{SO}(n)/\ker(\phi_{w})$ with the positive normalized $\operatorname{SO}(n)$-invariant measure $\sigma$ induced by $\widehat{\sigma}$ which exists and is unique. (See [29, Theorem 1 on p. 138] and use the fact that compact groups are unimodular for uniqueness.)
@@
$(\hat{\phi}_{w})^{-1}(g\Delta)$ $=\{g^{\prime}\in\operatorname{SO}(n)/\ker(\phi_{w})\colon g^{\prime}w\in g\Delta\}=\{g^{\prime}\in\operatorname{SO}(n)/\ker(\phi_{w})\colon g^{-1}g^{\prime}w\in\Delta\}$
@@
We already established the existence of a measure in Proposition 4.1. It remains to prove the uniqueness. Let us assume to the contrary that $\mu_{1}$ and $\mu_{2}$ are two different normalized $\operatorname{SO}(n)$-invariant measures on $T$. Then $(\phi^{-1}_{w})_{\ast}(\mu_{1})$ and $(\phi^{-1}_{w})_{\ast}(\mu_{2})$ are two different normalized $\operatorname{SO}(n)$-invariant measures on $\operatorname{SO}(n)/\ker(\phi_{w})$. But this contradicts the uniqueness of $\sigma$. ∎
@@
In [28, 11] the set $T$ is known as the Stiefel manifold $V_{2,n}(\mathbb{R})$ of all 22–frames in $\mathbb{R}^{n}$, i.e., the sets of all pairs of orthonormal vectors in $\mathbb{R}^{n}$. Equivalently, $V_{2,n}(\mathbb{R})$ is the set of all real $n\times 2$ matrices $X$ such that $X^{T}X$ is the $2\times 2$ identity matrix. Regarding $T$ as a manifold, it can also be equipped with the uniform normalized measure with respect to the action of the orthogonal group $\operatorname{SO}(n)$ [11, §1.4.3]. This measure coincides with the measure $\sigma$ introduced above.
@@ 4.2. Estimates
Now that we defined the measure on $T$, we can construct the appropriate sections of $\operatorname{Pos}_{\mathcal{Q}}$ and $\operatorname{Sos}_{\mathcal{Q}}$ and present the volume estimates for these sections (see Theorems 4.5, 4.6 below).
@@
[…]
@@
(4.2) $\mathcal{M}=\left\{f\in\mathcal{Q}\colon\int_{T}f\;{\rm d}\sigma=0\right\}.$
Since $\sigma$ is normalized,
@@
The proof of Lemma 4.4 is the same as the proof of [22, Lemma 1.4].
We are now ready to compare the volumes of the sections defined above. The lower bound for the volume of the section of nonnegative biforms from $\mathcal{Q}$ is as follows:
@@
For every $n\in\mathbb{N}$ the probability that a cross-positive map $\Phi:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ is completely cross-positive, is bounded above by
@@
Remark 4.9.
  1. (1)
    The correspondence (1.3) between $\ast$-linear maps $A:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ and biquadratic biforms $p_{A}$ is bijective only when maps are restricted to symmetric matrices $\mathbb{S}_{n}(\mathbb{R})$. Since nonnegativity of $p_{A}$ on $V(I)$ is equivalent to $A$ being cross-positive (see Proposition 2.5), while $p_{A}$ being a sum of squares modulo $I$ is equivalent to the existence of some completely cross-positive extension $\widetilde{A}:M_{n}(\mathbb{R})\to M_{n}(\mathbb{R})$ of the restriction of $A$ to $S_{n}(\mathbb{R})$ (see Proposition 2.6), $p_{n}$ is clearly an upper bound for the probability that a cross-positive map is completely cross-positive.
  2. (2)
    If we want to compare the sizes of two cones $K\subseteq L\subseteq\mathbb{R}^{n}$ in a fixed metric, then the most unbiased choice of a compact set $C$ to compare the sizes of $K\cap C$ and $L\cap C$ is the unit ball $B$ of this metric. In our case, the metric is the $L^{2}$ norm, coming from the inner product (4.3). In this norm, the condition $f\in B$ is given by a quadratic inequality in the coefficients of $f$ and therefore sharp lower and upper bounds on $K\cap B$ (resp. $L\cap B$) following the same asymptotics are difficult to establish. Replacing the unit ball $B$ with a hyperplane whose normal is some vector from the unit sphere leads to more manageable conditions. The choice of the hyperplane is not arbitrary, since its position can have a large impact on the size difference of the intersections, e.g., if the normal is almost perpendicular to some ray on the boundary of the larger cone $L$, then the difference in size can be very large, even if the smaller cone is not significantly smaller. However, if there is a vector in the interior of both cones, which is fixed by all symmetries for each cone, then the orthogonal complement of this vector is a fair choice of the hyperplane to capture size difference between the cones. In our case, the polynomial $(\sum_{i=1}^{n}x_{i}^{2})(\sum_{i=1}^{n}y_{i}^{2})$ is a fixed point for the action of the orthogonal group $O(n)$ on $\mathbb{R}[{\tt x},{\tt y}]_{2,2}$, defined by $O\cdot p({\tt x},{\tt y})=p(O^{-1}{\tt x},O^{-1}{\tt y})$. Note that the ideal $I$ is invariant under this action and therefore the action extends naturally to the action on $\mathcal{Q}$. It is clear that the sets $\operatorname{Pos}_{\mathcal{Q}}^{(n)}$ and $\operatorname{Sos}_{\mathcal{Q}}^{(n)}$ are invariant under this action and therefore comparing their sizes by intersecting them with $\mathcal{H}_{\mathcal{Q}}^{(n)}$ is an appropriate choice.
  3. (3)
    Blekherman [5, Theorem 6.1] established volume bounds for sum of squares forms. Our proofs of Theorems 4.5 and 4.6 freely borrows his ideas. An important ingredient in the proof of Theorem 4.6 is also a new version of the Reverse Hölder inequality, which we prove in Section 4.3 below.
  4. (4)
    In [5] Blekherman proved that for a fixed degree bigger than 2 the ratio between the volume radii of compact sections of the cones of sum of squares forms and nonnegative forms goes to 0, as the number of variables goes to infinity. Corollary 4.7 is an analog of his result for sum of squares biquadratic forms and nonnegative biquadratic forms on Stiefel manifolds $V_{2,n}(\mathbb{R}).$ (See Remark 4.3.)
Proof of Theorem 4.5.
We denote $\mathcal{K}=\widetilde{\operatorname{Pos}}^{(n)}_{\mathcal{Q}}$. As in [22, §2.1.1] we establish that
To prove this claim we will use [2, Corollary 2]. Let $(\mathbb{R}^{n}\otimes\mathbb{R}^{n})^{\otimes 2}$ be the $2$-nd tensor power of $\mathbb{R}^{n}\otimes\mathbb{R}^{n}$ . Let $e_{1},e_{2}\in\mathbb{R}^{n}$ be standard unit vectors and let $w$ be the tensor
The group $\operatorname{SO}(n)$ acts on $(\mathbb{R}^{n}\otimes\mathbb{R}^{n})^{\otimes 2}$ by the natural diagonal action, i.e., for $g\in\operatorname{SO}(n)$ and all $x_{i}\in\mathbb{R}^{n}$,
$g(x_{1}\otimes\cdots\otimes x_{4})=gx_{1}\otimes\cdots\otimes gx_{4}$
and extend by linearity. We also define
and we integrate with respect to the Haar measure $\widehat{\sigma}$ on $\operatorname{SO}(n)$. As in [3, Example 1.2], we proceed as follows:
  1. (1)
    Note also that $\operatorname{SO}(n)(e_{1}\otimes e_{2})=T$.
  2. (2)
  3. (3)
    We introduce an inner product on $V_{2}$ by defining
    This inner product also induces the dual inner product on the dual space $V_{2}^{\ast}\cong V_{2}$ which we also denote by $\langle\cdot,\cdot\rangle$.
By [2, Corollary 2],
where the equality follows as in [2, p. 404]. If $n$ is odd, we let $2k_{0}=9(n-1)$. Otherwise take $2k_{0}=9n$ to get
$D_{k_{0}}^{\frac{1}{2k_{0}}}\leq\binom{\frac{20}{9}k_{0}}{2k_{0}}^{\frac{1}{k_{0}}}.$
Since $2k_{0}=9\ell_{0}$ for some $\ell_{0}\in\mathbb{N}$,
where we used [22, Lemma 2.2] in the last inequality.
(4.4) $A=\int_{S_{\mathcal{M}}}\left\|f\right\|_{2k_{0}}\;{\rm d}\widetilde{\mu}=\int_{S_{\mathcal{M}}}\left(\int_{T}f^{2k_{0}}\;{\rm d}\sigma\right)^{\frac{1}{2k_{0}}}{\rm d}\widetilde{\mu}$.
Notice that
(4.5) $\int_{S_{\mathcal{M}}}\left(\int_{T}f^{2k_{0}}\;{\rm d}\sigma\right)^{\frac{1}{2k_{0}}}{\rm d}\widetilde{\mu}=\int_{c\in S_{V^{\ast}_{2}}}\left(\int_{g\in\operatorname{SO}(n)}\langle c,gv\rangle^{2k_{0}}{\rm d}\widehat{\sigma}(g)\right)^{\frac{1}{2k_{0}}}{\rm d}\breve{\sigma}(c),$
where $S_{V^{\ast}_{2}}$ is the unit sphere in $V^{\ast}_{2}$ endowed with the rotation invariant probability measure $\breve{\sigma}$. Combining (4.4), (4.5) we obtain
where we used [3, Lemma 3.5] for the inequality and [3, Remark p. 62] for the last equality. This equality proves the Claim and establishes the lower bound in Theorem 4.5. ∎
Proof of Theorem 4.6.
As in [22, §2.3.1], it follows that
$H_{f}(g)=\langle f,g^{2}\rangle\quad\text{for }g\in\mathcal{U}.$
Note that
$\|f\|_{\operatorname{sq}}=\|H_{f}\|_{\infty}.$
As in [22, p. 3343–3344] (for $k_{1}=k_{2}=1$), it follows that
where the last inequality follows by $\|g^{2}\|_{2}=\|g\|_{4}^{2}$ and Proposition 4.10 below. To prove the Claim it remains to establish
(4.6) $\sqrt{\frac{2D_{\mathcal{U}}}{D_{\mathcal{M}}}}\leq 2^{\frac{3}{2}}n^{-1}.$
The dimensions $_{\mathcal{U}}, DD_{\mathcal{M}}$ are easily verified to be
$D_{\mathcal{U}}$ $=$ $\dim\mathbb{R}[{\tt x},{\tt y}]_{1,1}-1=n^{2}-1,$
Observe that
$\underbrace{\leq}_{n>2}$ $\frac{2^{3}}{n^{2}+2n-2}\leq\frac{2^{3}}{n^{2}},$
which proves (4.6). ∎
@@ 4.3. Reverse Hölder inequality
We write $I_{1,1}=I\cap\mathbb{R}[{\tt x},{\tt y}]_{1,1}$. A bilinear form $g\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}/I_{1,1}$ is symmetric (resp. skew-symmetric) if it of the form $g({\tt x},{\tt y})={\tt x}^{T}A{\tt y}+I_{1,1}$ for some symmetric (resp. skew-symmetric) matrix $A\in\mathbb{R}^{n\times n}$.
@@
For a bilinear biform $g\in\mathbb{R}[{\tt x},{\tt y}]_{1,1}/I_{1,1}$,
@@
If $g$ is symmetric, then we can take $\sqrt{3}$ instead of $\sqrt{6}$ in (4.7) above, while if $g$ is skew-symmetric, $\sqrt{6}$ can be replaced by $\sqrt[4]{6}$. Moreover, the constants $\sqrt{3}$ (resp. $\sqrt[4]{6}$) are asymptotically sharp as $n\to\infty$.
@@
The proof of Proposition 4.10 will be done separately for the symmetric (Section 4.3.2) and skew-symmetric case (Section 4.3.3), while the general case (Section 4.3.4) follows from the fact that every bilinear form $g$ can be written as a sum of a symmetric form $g_{s}$ and a skew-symmetric form $g_{a}$, together with the observation that $g_{s}$ is perpendicular to $g_{a}$ in the $L^{2}$ inner product. For the proof we first need to compute the values of the integrals of monomials of bidegree $(2,2)$ and some monomials of bidegree $(4,4)$ with respect to $\sigma$, which is the content of Section 4.3.1. Using these computations, (4.7) becomes an inequality in the coefficients of $g_{a}$ (resp. $g_{s}$). We prove that this inequality holds.
@@
In [16] a version of the Reverse Hölder inequality with respect to the Lebesgue measure on the unit sphere and polynomials of any degree is established. Lemma 2.9 of [22] extends this result to the product measure of two Lebesgue measures on unit spheres. However, in the proof of Theorem 4.6 we cannot use this extension because the measure $\sigma$ is not a product measure. Therefore we have to establish the Reverse Hölder inequality we need in our setting.
@@
[…]
@@
$\text{for }i,j,k,l=1,\ldots,n,\;i,j,k\text{ pairwise different},$
@@
$\text{for }n\geq 4\text{ and }i,j,k=1,\ldots,n,\;i,j,k,l\text{ pairwise different},$
@@
$\text{for }n\geq 4\text{ and }i,j,k,l\text{ pairwise different},$
@@
[…]
@@
[…]
@@
Then $H_{n}^{1}(\underline{\phi})$ is (see, e.g., the formula for $L_{1}(\theta)$ in [42, p. 3–4] or use induction on $n$)
@@
The set $T=V_{2,n}(\mathbb{R})$ (see Remark 4.3) can be parametrized by (see, e.g., [11, p. 48–49] or [42, §2])
@@
where by [10, Theorem 2.1] (taking $V=x(\underline{\phi})$, $G(V)=H_{n}^{1}(\underline{\phi})\text{ without the first column}$ and $Z=\text{the first column of }H_{n-1}^{1}(\underline{\psi})$) and [7],
@@
By the invariance of the integral with respect to the change of indices we can assume without loss of generality that $i,j,k,l\in\{1,2,3,4\}$ in all equalities of Lemma 4.12. Due to the difference in parameterizations of some of the coordinates $x_{i}(\underline{\phi})$, $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3,4$, we separate cases $n\geq 6$, $n=5$, $n=4$ and $n=3$ in the rest of the proof.
Case 1: $n\geq 6$. The coordinates $x_{i}(\underline{\phi})$, $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3,4$, are the following:
@@
The fact $I_{3}=\frac{n+1}{n-1}I_{1}$ is a consequence of the following computation:
$1$ $=\int_{T}\Big{(}\sum_{i=1}^{n}{\tt x}_{i}^{2}\Big{)}\Big{(}\sum_{i=1}^{n}{\tt y}_{i}^{2}\Big{)}\;{\rm d}\sigma=\int_{T}\Big{(}\sum_{i=1}^{n}{\tt z}_{i}^{2}\Big{)}\;{\rm d}\sigma+\int_{T}\Big{(}\sum_{i\neq j}{\tt z}_{ij}^{2}\Big{)}\;{\rm d}\sigma=nI_{1}+n(n-1)I_{3}.$
@@
Next we prove that $\int_{T}{\tt z}_{ij}{\tt z}_{kl}\;{\rm d}\sigma=0$ if at least one of $i,j,k,l$ occurs an odd number of times. Write $g({\tt x},{\tt y}):={\tt x}_{i}{\tt y}_{j}{\tt x}_{k}{\tt y}_{l}$ and let $i_{1}$ be the index among $i,j,k,l$ which occurs an odd number of times. Since
@@
Case 2: $n=5$. Note that the parameterizations of $x_{i}(\underline{\phi})$, $i=1,2,3,4$, and $y_{i}(\underline{\phi},\underline{\psi})$, $i=1,2,3$, are the same as in the case $n\geq 6$, while
@@
$0$ $=\int_{T}({\tt z}_{1}+{\tt z}_{2}+{\tt z}_{3}+{\tt z}_{4}+{\tt z}_{5})^{4}\;{\rm d}\sigma=\sum_{i=1}^{5}\int_{T}{\tt z}_{i}^{4}\;{\rm d}\sigma+4\sum_{i\neq j}\int_{T}{\tt z}_{i}^{3}{\tt z}_{j}\;{\rm d}\sigma+6\sum_{i<j}\int_{T}{\tt z}_{i}^{2}{\tt z}_{j}^{2}\;{\rm d}\sigma$